process log status of working monitor data during ep rinsing process comparison of ep2 n.
Download
Skip this Video
Loading SlideShow in 5 Seconds..
Process log Status of working Monitor data during EP & Rinsing process Comparison of EP2 PowerPoint Presentation
Download Presentation
Process log Status of working Monitor data during EP & Rinsing process Comparison of EP2

Loading in 2 Seconds...

play fullscreen
1 / 20

Process log Status of working Monitor data during EP & Rinsing process Comparison of EP2 - PowerPoint PPT Presentation


  • 121 Views
  • Uploaded on

MHI#14 号機の2回目の EP1. Period : 2011/4/19 ~ 4/21 Process : EP1(100 μ m, 50mA/cm 2 ), Normal 1 st Water Rinsing [90min], Degreasing (FM-20, 2%) [30min], HPR [6 hours] Preparation :Attachment to jigs Workers : M. Sawabe, Kirk, M. Satoh (KEK),

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about 'Process log Status of working Monitor data during EP & Rinsing process Comparison of EP2' - harding-mercer


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
process log status of working monitor data during ep rinsing process comparison of ep2

MHI#14号機の2回目のEP1

Period : 2011/4/19 ~4/21

Process : EP1(100μm, 50mA/cm2), Normal 1st Water Rinsing[90min],

Degreasing (FM-20, 2%) [30min], HPR [6 hours]

Preparation :Attachment to jigs

Workers : M. Sawabe, Kirk, M. Satoh (KEK),

K. Nakamura, N. Tasaki, F. Tsukada (Assist Engineering Co.),

T. Okada (K-VAC),

M. Asano, S. Imada, T. Yanagimachi (Nihon Advanced Technology)

Process log

Status of working

Monitor data during EP & Rinsing process

Comparison of EP2

STF Cavity Group Meeting @2011/4/25

Kirk

condition of ep rinsing
Condition of EP & Rinsing
  • EP acid (4/20のEP1後のサンプルを使用)
      • Nb ingredient : 9.6g/ℓ
      • HF ingredient : ? g/ℓ (全体?g/ℓの65%が反応に使われる)
      • Al ingredient : 18mg/ℓ
      • Current density : ~50mA/cm2 (EP1の時は通常の電流密度で行う)
      • Cavity surface temperature : <40℃
      • Control Voltage : 温度と電流密度を見ながら制御
      • With normal N2 gas flow during extracting EP acid (8ℓ/min)
  • 1st water rinsing (1.5hour)
      • 10min (storing)/3min (flushing)で行う
  • Degreasing (30min)
      • FM-20 (2%)
  • HPR (~6 hours)
      • 28m (1 turn, w/o blind flanges)
      • 5h16m (6 turns, w/o blind flanges)

STF Cavity Group Meeting @2011/4/25

process log
Process log ①

4/19 9:00 MHI#14 is moved to EP area

10:00 Check every flange [O.K.] (polishing HOM at input side)

10:05 Attachment of jigs for EP

10:30 HPR starts without cavity

11:00 Meeting for tomorrow’s working procedure

11:20 Set cavity to EP bed with crane

11:25 Installation of cathode bar to cavity

11:40 Set data logger to cavity (totally 12ch)

Leak check with N2 gas [O.K.]

13:15 fin.

STF Cavity Group Meeting @2011/4/25

process log1
Process log ②

4/20 8:35 Cavity rotation starts / EP acid flow starts

8:59 EP1 starts

10:38 Alarm occurred due to low acid level in reservoir! @36.6μm

Stop current flow and acid flow

It was not recover even after pushing “Reset” button many times

Check acid level meter in reservoir (dangerous working!)

11:02 EP acid flow restarts

11:04 EP1 restarts (Set 65μm)

13:58 EP1 stops [5h00m] / Idling rotation (3 rpm / 20min) / Plastic case detachment

14:16 Idling rotation stops

14:18 EP acid is removed from cavity with N2 gas flow(0.020MPa)

14:32 First rinsing with ultra pure water starts

16:02 First rinsing with ultra pure water stops [1h30m]

16:11 Detachment of restriction jigs / Data logger detached

16:25 Un-installation of cathode bar from cavity / washing cathode bar (No black mark)

16:35 transferring cavity from EP bed to wagon

16:58 Detachment of jigs for EP / Detaching every blind flange of teflon

/ Checking inside cavity (No mark) / Checking teflon surface (No mark)

17:06 Shower rinsing @CP area

17:15 Attachment of flanges for degreasing

17:20 Supplying degreaser into cavity

17:27 Degreasing with ultrasonic starts

17:57 Degreasing with ultrasonic stops [30min]

STF Cavity Group Meeting @2011/4/25

process log2
Process log ③

4/20 18:00 Detachment of flanges for degreasing

18:15 Shower rinsing @CP area

18:25 Transferring cavity from wagon on turntable for HPR using crane

18:38 Checking tolerance between position of nozzle and center of beam pipe

18:40 1st HPR for inside cavity starts

19:08 1st HPR for inside cavity stops [0h28]

19:10 fin

4/21 9:08 2nd HPR starts

14:25 2nd HPR stops [5h17m] / Drying @HPR room

4/22 9:00 move MHI#14 to inspection room

4/25 9:00 optical inspection starts

STF Cavity Group Meeting @2011/4/25

status of working
Status of working ①

STF Cavity Group Meeting @2011/4/25

status of working1
Status of working ②

STF Cavity Group Meeting @2011/4/25

status of working2
Status of working ③

STF Cavity Group Meeting @2011/4/25

status of working3
Status of working ④

STF Cavity Group Meeting @2011/4/25

status of working4
Status of working ⑤

STF Cavity Group Meeting @2011/4/25

status of working5
Status of working ⑥

STF Cavity Group Meeting @2011/4/25

status of 2 nd ep1 for mhi 14
Status of 2nd EP1 for MHI#14 ①

current

Tcavity

Treservoir

voltage

Troom

average current density

EP1の場合は通常通り50mA/cm2で行う。

途中で2回ほどジャンプしているのは、

電圧を20V前後で調整していたことによる。

また、リザーバタンク内の液面計のトラブルに

因り、20分ほどの中断があった。

Tcavity

STF Cavity Group Meeting @2011/4/25

status of 2 nd ep1 for mhi 141
Status of 2nd EP1 for MHI#14 ②

左図は平均電流密度と電圧の相関

右図は定常状態での分布

STF Cavity Group Meeting @2011/4/25

slide14
最近のEP1の定常状態の比較

MHI#17のEP1の結果はドリフトの効果が大きい。

STF Cavity Group Meeting @2011/4/25

status of 2 nd ep1 for mhi 142
Status of 2nd EP1 for MHI#14 ③

EP1の間の空洞表面温度の状況。

テープが取れてしまった

#1

#2

#3

#4

#5

#6

#7

#8

#9

#10

#11

#12

STF Cavity Group Meeting @2011/4/25

Steady state: 38.5 35.1 34.5 32.4 34.3 20.4 30.8 33.7 32.3 37.0 35.4 11.3

status of 2 nd ep1 for mhi 143
Status of 2nd EP1 for MHI#14 ④

EP1の間の空洞表面温度の状況(前ページの最後の拡大図)。

テープが取れてしまった

#1

#2

#3

#4

#5

#6

#7

#8

#9

#10

#11

#12

STF Cavity Group Meeting @2011/4/25

status of uwr for mhi 14
Status of UWR for MHI#14 ①

#1

#2

#3

#4

#5

#6

#7

#8

#9

#10

#11

#12

STF Cavity Group Meeting @2011/4/25

Steady state: 19.1 18.2 18.8 17.0 17.9 17.5 17.3 18.5 17.7 18.2 18.3 17.7

status of uwr for mhi 141
Status of UWR for MHI#14 ②

一次洗浄中の空洞温度の状況(前ページの拡大図)

#1

#2

#3

#4

#5

#6

#7

#8

#9

#10

#11

#12

STF Cavity Group Meeting @2011/4/25

status of uwr for mhi 142
Status of UWR for MHI#14 ③

STF Cavity Group Meeting @2011/4/25

status of hpr for mhi 14
Status of HPR for MHI#14

クリーンルーム作業を始めるために、クラス1000のフィルターのファンを入れたら、TOCが徐々に上がっていった。

STF Cavity Group Meeting @2011/4/25